OpenStudy (anonymous):

Chapter 5:Indices and Logarithms @mathmath333

OpenStudy (anonymous):

\[Given~\log _{x}p=3~and~\log _{x}q=5,\]find the value of:\[\log _{x}\frac{ \sqrt{p} }{ q }\]

OpenStudy (anonymous):

\[p=x^3\]\[q=x^5\]\[\log _{x}\frac{ \sqrt{p} }{ q }=\log_{x}\frac{x^{3/2}}{x^5}=\log_{x}x^{\frac{ 3 }{ 2 }-\frac{ 10 }{ 2 }}=\log_{x} x^{\frac{-7}{2}}=\frac{-7}{2}\log_{x} x=\frac{-7}{2}\]

OpenStudy (anonymous):

Thnx @Nishant_Garg

OpenStudy (anonymous):



OpenStudy (anonymous):

??